assume a random sample of n measurements is selected from population with sample size of 25 and standard deviation of 40?whats the 0-x. a.20b.16c.8d.10
Answer #1

Let X1,X2,•••••••Xn are random sample of size n taken from population

Such that

random sample size =n=25

and standard deviation of X=40

we have to calculate standard deviation of

First let's calculate variance of =Var ((X1+X2+•••+Xn)/n)=Var(X1+X2+•••+Xn)/n^2

=n×Var(X)/n^2=Var(X)/n. {due to all random sample are coming from same population with variance is 40^2}

=40^2/25=1600/25=64

Standard deviation of =sqrt(64)=8 Answer is (c)